APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Wednesday October 12, 2005 (MJD 53656) ---=== OBSERVING TEAM ===--- Swing: John Barentine Night: Jurek Krzesinksi Howard Brewington superNova: Ben Dilday Training: Viktor Malanushenko Olena Malanushenko Support: Eric, Kurt, and Connie ---=== OBSERVING PLAN ===--- Shakeup and SN Science ---=== OBSERVING SUMMARY ===--- We did 82N (superNova) 115 to 176 (excluding ramps) and, we spooled the data. We completed the imager monthly checkout. directory for bias/flat check sdsshost2:/home/observer/bias/biascheck53656 Closed PR 6627, poolCopy change Closed PR 6635, DA Transfer PR6614 and PR6645 were not closed; see log. PR6636 was not closed either; we sent Kurt an emial about this task. iop was not declared current; it needs more testing DIMM was powered down (thanks Steph) No runLogs created after superNova Run: 5719. We tried manually but had not success. Eric was notified The imagingSummary has five 5719 runs because of endDrift problems. ---=== OBSERVING LOG ===--- Afternoon --------- Setup iop and astroDa to run test versions (4_6_0 and 15_5, respectively). See Problems section for details. No problems during checkout. Night ----- With Ben's help (superNova guy), we mounted the imager and pulled the enclosure at sunset. We plan to do superNova during the first part of the night and shake during the second. So, we moved to 82N -startLambda 115 and started a drift. During the superNova run, we got a Watcher message that id-1 was not being archived; see message below. dsc E archWrFail Fatal error or retries exhausted. File idRR-005719-g1-0284.fit not archived! (dsc_E_archWrFail Fatal error or retries exhausted. File idRR-005719-g1-0284.fit not archived!) Next, Watcher gave us Micro/imager message; see message below. server is not responding (Last packet: 2005-10-13 04:17:39Z) We called Connine, and she asked that we do a cc camcheck, which reported no problems. Then, the servers crashed, and there was also a VME problem. And, we could not endDrift since our iop sessions were crashing too. We called Eric, and he suggested that we rebooted the crates. Rebooting the crates did fix the problem however, endDrift was a little confused, i.e., the frames were misnumbered; see message below. And there was no runLog created either. We tried the makeRunLog command twice but realized no success; see below. im> makeRunLog [mjd4Gang] 5719 filling out runlog /astrolog/53656/runlog-5719.txt for run 5719. Error: can't read "noise": no such variable im> --------- im> endDrift 2005-10-13 06:15:29Z: stopping drift with: iccExec id1 id2 id3 id4 id5 id6 ad1 ad2 ad3 ad4 readoutEnd 2005-10-13 06:15:59Z: endData has failed: {id1 {ICC id1 (crate 1, board 0) is not available Bad ICC}} {id2 {ICC id2 (crate 1, board 1) is not available Bad ICC}} {id3 {ICC id3 (crate 1, board 2) is not available Bad ICC}} {id4 {}} {id5 {}} {id6 {}} {ad1 {ICC ad1 (crate 3, board 0) is not available Bad ICC}} {ad2 {ICC ad2 (crate 3, board 1) is not available Bad ICC}} {ad3 {ICC ad3 (crate 3, board 2) is not available Bad ICC}} {ad4 {ICC ad4 (crate 3, board 3) is not available Bad ICC}} Flavor of this run (science, bias, calibration, engineering, ignore, help) [science]: 82,N Invalid flavor 82,N. Please try again Flavor of this run (science, bias, calibration, engineering, ignore, help) []: science Stripe, Strip (e.g, 22,N or SS,N) []: 82,N Mountain estimate of the overall data quality (unknown, bad, acceptable, good) [unknown]: Please provide a one line comment [] leap seconds = 32 Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 5716 21:29Z 21:52Z 100 O 53.08 58.79 51 ignore 5717 22:37Z 00:00Z 100 O 70.11 90.93 12 ignore 5719 01:38Z 06:17Z 82 N 111.33 -178.74 -1 science First good frame of this run (8): First good frame of this run (8): 456 First good frame of this run (8): First good frame of this run (8): First good frame of this run (8): First good frame of this run (8): First good frame of this run (8): First good frame of this run (8): 8 First good frame of this run (8): 1 First good frame of this run (8): 10 First good frame of this run (8): quitting if breakCheck called..., or tell me three times (1) Error: error reading "stdin": Interrupted system call We ctrl+c that and used Eric's fix: im> source /home/neilsen/iop/etc/highCommand.tcl im> endDrift worked. ----== Shake Items ==---- The following marked, **(YES)**, worked. And the ones marked **(NO)** did not work. # Imaging bias run 1. verify that the ramp-up dummy frames all show up in /data/photo/frames within the first minute or so of the run. **(YES)** 2. verify that data frames are copied to /data/photo/frames as soon as they are generated. **(YES)** 3. verify that the ramp-down dummy frames all show up on the Linux disk soon after the end of the run. **(YES)** 4. look at an uncompressed photo frame file (using more or a similar tool) to verify that the DATASUM and CHECKSUM keywords are in the header.**(YES)** 5. verify that the FITS header checksum is correct for a few files. To do this, you should catch the files before they are compressed, run ~products/.biery/fitsChecksum , and verify that the output is "checksum value is 0xffffffff, total characters = 5803200".**(YES)** 6. check that frameCatcher is correctly finding the FITS header DATASUM and using it to validate the integrity of the frame before compressing it..**(YES)** 7. in a long run, check for occasional FTP failures. These should show up as a murmur warning message (dsc_W_archWrAttempt) followed by a retry message (dsc_W_archWrRetry) followed by either a success message (TEXTONLY write filename ) or a failure message (dsc_E_archWrFail). **(YES)** # Spectro bias run 1. verify that the frames (once they are copied to the Linux disk) have the FITS header DATASUM and CHECKSUM keywords and that the checksum is correct. **(NO)** 2. verify that the permissions on the new files in /data/spectro are 0664. **(YES)** # PT run 1. verify that the frames (once they are copied to the Linux disk) have the FITS header DATASUM and CHECKSUM keywords and that the checksum is correct. **(NO)** # Use poolCopy to transfer some files by hand and verify that the permissions on the files are all equal to 0664.**(YES)** # Use dsclat to connect to each node (id1 - id6, ad1 - ad4, sd1, md1). Verify with a that the correct node is there. *** We connected to all the nodes and found the correct node each time, but had trouble connecting to id1 was busy. We talked to Kurt about this, and he said that he would talk to Jon Brinkmann on Thursday morning. _______________________________________ We also tested poolCopy: first we checked what is in the pool: >poolDir ... 5713 621.41 --c-a-- 5792416 12-Oct-2005 07:30:27.76 TAI 5713 621.51 --c-a-- 5792416 12-Oct-2005 07:30:27.76 TAI 5713 622.11 --c-a-- 5792416 12-Oct-2005 07:31:03.59 TAI 5713 622.21 --c-a-- 5792416 12-Oct-2005 07:31:03.59 TAI 5714 12.0 -gcra-- 388800 12-Oct-2005 08:52:50.87 TAI then im> poolCopy 5714 12.0 ala.fits and the file has permisions set like this: -rw-rw-r-- 1 observer dss 397440 Oct 13 07:13 ala.fits So the test has passed, but when we looked at the fits file header for that file: im> more ala.fits SIMPLE = F / FIT STANDARD BITPIX = 8 / FITS BITS/PIXEL NAXIS = 0 / No axes EXTEND = T / Fits extensions are used SDSS / THIS HEADER CONTAINS SPECIAL SDSS KEYWORDS NAXIS is 0 -is it right? Imager monthly checkout --------------------------- run 5721 - bias on the telescope >getAllFrames 5721 35 . im> checkBias 5721 35 . {61 WARN: } {61 FAIL: leftRNComp } <==== problems! Open /skycat/imager/53656/5721-35-u-bias.gif/gif Open /skycat/imager/53656/5721-35-g-bias.gif/gif Open /skycat/imager/53656/5721-35-r-bias.gif/gif Open /skycat/imager/53656/5721-35-i-bias.gif/gif Open /skycat/imager/53656/5721-35-z-bias.gif/gif Open /skycat/imager/53656/5721-35-o-bias.gif/gif Open /skycat/imager/53656/5721-35-l-bias.gif/gif Open /skycat/imager/53656/5721-35-t-bias.gif/gif Open /skycat/imager/53656/5721-35-s-bias.gif/gif {photo h1} {astro h2} --------------------------------------------------- im> biasGen 5721 Directory /data/photo/53656/5721, Field 40 Writing files to directory /skycat/bias/53656 Camcol 1 Row 1 Camcol 1 Row 2 Camcol 1 Row 3 Camcol 1 Row 4 Camcol 1 Row 5 Camcol 2 Row 1 Camcol 2 Row 2 Camcol 2 Row 3 Camcol 2 Row 4 Camcol 2 Row 5 Camcol 3 Row 1 Camcol 3 Row 2 Camcol 3 Row 3 Camcol 3 Row 4 Camcol 3 Row 5 Camcol 4 Row 1 Camcol 4 Row 2 Camcol 4 Row 3 Camcol 4 Row 4 Camcol 4 Row 5 Camcol 5 Row 1 Camcol 5 Row 2 Camcol 5 Row 3 Camcol 5 Row 4 Camcol 5 Row 5 Camcol 6 Row 1 Camcol 6 Row 2 Camcol 6 Row 3 Camcol 6 Row 4 Camcol 6 Row 5 Rank 6 CCD 0 Rank 6 CCD 1 Rank 6 CCD 2 Rank 6 CCD 3 Rank 6 CCD 4 Rank 6 CCD 5 Rank 7 CCD 1 Rank 7 CCD 2 Rank 7 CCD 3 Rank 7 CCD 4 Rank 7 CCD 5 Rank 7 CCD 6 Rank 8 CCD 1 Rank 8 CCD 2 Rank 8 CCD 3 Rank 8 CCD 4 Rank 8 CCD 5 Rank 8 CCD 6 Rank 9 CCD 0 Rank 9 CCD 1 Rank 9 CCD 2 Rank 9 CCD 3 Rank 9 CCD 4 Rank 9 CCD 5 im> -------------------------------------------------------- im> biasCheck 5721 Directory /data/photo/53656/5721, Field 40 Select reference mjd 0. 53418 1. 53606 2. 53656 Select index of reference mjd [2]: 1 Writing file /astrolog/53656/idBiasCheck-005721.par CCD 11 Offset. Left: -9.02 +/- 9.97 Right: -9.03 +/- 9.98 CCD 21 Offset. Left: 3.00 +/- 0.00 Right: 4.50 +/- 0.53 CCD 31 Offset. Left: -0.19 +/- 0.58 Right: -0.90 +/- 0.42 CCD 41 Offset. Left: 5.00 +/- 0.00 Right: 5.17 +/- 0.38 CCD 51 Offset. Left: -28.50 +/- 12.46 Right: -28.60 +/- 12.39 CCD 12 Offset. Left: 3.00 +/- 0.00 Right: 5.00 +/- 0.00 CCD 22 Offset. Left: 4.00 +/- 0.00 Right: 8.34 +/- 0.48 CCD 32 Offset. Left: -71.23 +/- 0.59 Right: 1.00 +/- 0.00 CCD 42 Offset. Left: 4.75 +/- 0.43 Right: 5.00 +/- 0.00 CCD 52 Offset. Left: 1.00 +/- 0.00 Right: 0.00 +/- 0.00 CCD 13 Offset. Left: 2.75 +/- 0.44 Right: 2.26 +/- 0.44 CCD 23 Offset. Left: 0.95 +/- 0.65 Right: 3.00 +/- 0.00 CCD 33 Offset. Left: 0.54 +/- 0.53 Right: -3.90 +/- 0.30 CCD 43 Offset. Left: -0.19 +/- 0.39 Right: -1.00 +/- 0.00 CCD 53 Offset. Left: -3.26 +/- 0.44 Right: 3.00 +/- 0.00 CCD 14 Offset. Left: 4.00 +/- 0.00 Right: 4.00 +/- 0.00 CCD 24 Offset. Left: 5.00 +/- 0.00 Right: 6.00 +/- 0.00 CCD 34 Offset. Left: -34.46 +/- 2.63 Right: -34.16 +/- 2.86 CCD 44 Offset. Left: 3.00 +/- 0.00 Right: 1.10 +/- 0.31 CCD 54 Offset. Left: 3.14 +/- 0.35 Right: 7.65 +/- 0.48 CCD 15 Offset. Left: 4.00 +/- 0.00 Right: 5.00 +/- 0.00 CCD 25 Offset. Left: 12.65 +/- 0.53 Right: 12.69 +/- 0.46 CCD 35 Offset. Left: -5.17 +/- 4.88 Right: -1.70 +/- 0.48 CCD 45 Offset. Left: 2.00 +/- 0.00 Right: 2.00 +/- 0.00 CCD 55 Offset. Left: 3.00 +/- 0.00 Right: 4.00 +/- 0.00 CCD 16 Offset. Left: 8.00 +/- 0.00 Right: 9.30 +/- 0.48 CCD 26 Offset. Left: 1.23 +/- 0.43 Right: 1.67 +/- 0.47 CCD 36 Offset. Left: -18.58 +/- 0.51 Right: -18.72 +/- 0.49 CCD 46 Offset. Left: 1.00 +/- 0.00 Right: 3.66 +/- 0.50 CCD 56 Offset. Left: 1.00 +/- 0.00 Right: 6.00 +/- 0.00 CCD 60 Offset. Left: 35.89 +/- 0.40 Right: 35.42 +/- 0.55 CCD 61 Offset. Left: 14.29 +/- 0.45 Right: 14.55 +/- 0.50 CCD 62 Offset. Left: 39.26 +/- 0.52 Right: 39.24 +/- 0.43 CCD 63 Offset. Left: 17.00 +/- 0.36 Right: 16.70 +/- 0.52 CCD 64 Offset. Left: 30.54 +/- 0.57 Right: 30.58 +/- 0.58 CCD 65 Offset. Left: -6.90 +/- 1.05 Right: -6.77 +/- 0.86 CCD 71 Offset. Left: 61.88 +/- 0.34 Right: 62.00 +/- 0.00 CCD 72 Offset. Left: 29.79 +/- 0.41 Right: 29.89 +/- 0.32 CCD 73 Offset. Left: -17.09 +/- 0.36 Right: -17.29 +/- 0.52 CCD 74 Offset. Left: 50.98 +/- 0.57 Right: 51.08 +/- 0.26 CCD 75 Offset. Left: 5.17 +/- 0.37 Right: 5.41 +/- 0.49 CCD 76 Offset. Left: 8.58 +/- 0.49 Right: -83.00 +/- 223.03 <=== CCD 81 Offset. Left: 28.56 +/- 0.51 Right: 28.42 +/- 0.99 CCD 82 Offset. Left: 38.48 +/- 0.98 Right: 38.75 +/- 0.47 CCD 83 Offset. Left: 37.90 +/- 0.31 Right: 37.87 +/- 0.60 CCD 84 Offset. Left: 16.57 +/- 0.51 Right: 16.71 +/- 0.48 CCD 85 Offset. Left: 16.46 +/- 0.51 Right: 16.47 +/- 0.50 CCD 86 Offset. Left: 20.10 +/- 0.30 Right: 20.14 +/- 0.35 CCD 90 Offset. Left: 18.93 +/- 0.40 Right: 18.87 +/- 0.43 CCD 91 Offset. Left: 15.89 +/- 0.35 Right: 16.26 +/- 0.44 CCD 92 Offset. Left: 17.95 +/- 0.31 Right: 17.77 +/- 0.48 CCD 93 Offset. Left: 32.85 +/- 0.44 Right: 32.93 +/- 0.38 CCD 94 Offset. Left: 21.63 +/- 0.48 Right: 21.62 +/- 0.49 CCD 95 Offset. Left: 38.87 +/- 0.33 Right: 38.99 +/- 0.15 im> run 5722: UVWht lamps on im> getAllFrames 5722 35 im> checkFlat 5722 35 . {42 WARN: leftGain } {42 FAIL: } Open /skycat/imager/53656/5722-35-u-flat.gif/gif Open /skycat/imager/53656/5722-35-g-flat.gif/gif Open /skycat/imager/53656/5722-35-r-flat.gif/gif Open /skycat/imager/53656/5722-35-i-flat.gif/gif Open /skycat/imager/53656/5722-35-z-flat.gif/gif 0 im> endNight finished and was happy. We also spooled the superNova data. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 5719 01:38Z 06:17Z 82 N 111.33 -178.74 465 science 5719 01:38Z 06:39Z 82 N 111.33 -64.00 -1 science ignore 5719 01:38Z 06:17Z 82 N 111.33 -178.74 -1 science ignore 5719 01:38Z 06:39Z 82 N 111.33 -64.00 -1 science ignore 5719 01:38Z 06:42Z 82 N 111.33 -64.00 -1 science ignore 5720 06:46Z 06:54Z 100 O -167.36 -165.40 27 ignore 5721 06:55Z 07:16Z 100 O -165.09 -159.85 49 bias 5722 07:50Z 08:09Z 100 O -151.23 -146.50 46 calibration 5723 08:56Z 09:07Z 100 O -134.86 -132.08 33 ignore ---=== IMAGING RUN DETAILS ===--- Run: 5718 ignore +++++++++++++++++++++++++ Run: 5719 superNova Science Started the run ~025 min before 18 deg twilight to get set up and hit the ramp at 18 deg twilight. Seeing quality as reported by seeingRun was about 1.8 to 2.0" Frame Lambda Notes ----- ------ -------------------------------------------------------- 09 111.5 Focus good from start 12 112.0 Skippy happy 18 113 Start ramp 32 115 Start science 442 174 End Science 456 176 End ramp ++++++++++++++++++++++++ Run: 5721 On-telescope bias ++++++++++++++++++++++++ ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- 5719 12 3 8 -1.8 0.567 -2.3 1.885 .0002 2.8 -2.5 5719 52 3 100 -5.3 0.171 -3.1 0.165 -.002 9.0 -3.7 5719 102 3 86 -5.4 0.169 -2.0 0.174 -.002 9.5 -2.6 5719 152 3 121 -5.0 0.153 -0.9 0.205 -.002 9.0 -1.4 ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 06:56 34242 0.0 bias 07:05 34244 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 21:06Z 55 30 11 231 (SW) -9999 - 06:19Z 41 32 19 84 (E) -9999 - 06:50Z 40 31 21 76 (ENE) -9999 - 07:20Z 40 31 22 73 (ENE) -9999 - 07:52Z 40 31 21 72 (ENE) -9999 - 08:24Z 40 31 21 73 (ENE) -9999 - 08:57Z 40 31 17 75 (ENE) -9999 - 09:30Z 41 31 21 74 (ENE) -9999 - 10:02Z 41 31 18 76 (ENE) -9999 - 10:34Z 41 31 12 87 (E) -9999 - 11:07Z 40 31 17 80 (E) -9999 - 11:39Z 40 31 23 80 (E) -9999 - ---=== TELESCOPE STATUS ===--- 00:05Z Cracked doors, opened louvers, fans on. 00:45Z Enclosure off 06:45Z Enclosure on, fans off, doors and louvers closed 08:59Z Doors cracked, fans on, louvers opened 11:04Z Doors closed, louvers closed, fans off Telescope at 121,30,00 mount LN2 connected and turned on Cart 05 mounted; CW at 240lbs Imager 198lbs 21psi Spectro 172lbs 24psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_6_0 Watcher: v2_30_2 MCP: v5_26_0 TPM: tpm_v3_1_0 AstroDa: v15_5 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.8400 -5.8440 Axial B -5.6690 -5.6770 Axial C 0.9360 0.9420 Trans D -9.9890 -9.9940 Lateral E 10.3759 10.3759 Lateral F 11.9507 11.9380 GALILS Commanded: 4700. -2450. 3000. -7800. 3300. 3000. Actual: 4711. -2457. 2997. -7825. 3296. 3024. SETMIR VALUES PriDesOrient: 0.00 -1.68 7.05 385.25 92.02 PriOrient: 0.00 -1.52 7.02 384.99 93.81 SECONDARY: ---------- Focus: 0.00 Air Temp.: 5.0 Alt.: 30.000708 MIGS TONIGHT NOMINAL Axial A 1.5470 1.5420 Axial B 1.0540 1.0470 Axial C 1.1880 1.1650 Trans D -0.1080 -0.1240 GALILS Commanded: -194084. -184352. -225883. -7100. -6000. Actual: 1593447. 1603642. 1554316. -6820. -5882. SETMIR VALUES SecDesOrient: 1257.00 15.87 0.00 -50.00 136.76 SecOrient: -160.04 13.23 0.38 -49.93 138.23 ---=== PROBLEMS IN DETAIL ===--- ---== Misc Watcher Messages ==--- m1 temp (Excessive Telescope - Ambient Difference:Excessive Front-Back Difference (Right Half):Excessive Front-Back Difference (Upper Half): 2005-10-13 00:57:13Z ) env temp (Excessive Return - Ambient Difference:Excessive Floor - Ambient Difference: 2005-10-13 00:57:13Z) tmicro (Disconnected Sensor in M2B4:: 2005-10-13 00:01:19Z) tracking (Excessive Az Tracking Error:: 2005-10-13 00:58:35Z) Text="failed to move one or more axes":AxisNoTrack:"TTT" (2005-10-13 00:32:57Z) Modu="axe i SchMove" (2005-10-13 00:33:24Z) Modu="axe o Init" (2005-10-13 00:33:24Z) Modu="exe Track" (2005-10-13 00:33:24Z) Modu="prt ReadReplySet" (2005-10-13 00:32:57Z) missing files: idRR-005719-g1-0276.fit idRR-005719-u1-0276.fit idRR-005719-i1-0284.fit idRR-005719-z1-0283.fit idRR-005719-z1-0288.fit idRR-005719-r1-0272.fit but qick check on flies1 showed that we've got them! PR was filed --------------------------------------- Well, I Do Declare... --------------------- As part of tonight's shakeup requests, we're testing new versions of IOP (v4_6_0) and astroDa (v15_5). I spoke to both Steph and Eric by phone who suggested setting up the test versions and then starting IOP as usual. Trying this I got the error message: Switching umask from 0022 to 002 warning: no nodes are configure to boot system v15_5 correct by using the -s with -f option to one of the following systems: OR by using iccConfig to configure a new system OR by setting up a different version of astroda Eric suggested the DA nodes would have to be pointed to the new version separate from the *OPs prior to starting up an *OP Checking Kurt's notes regarding the new astroDa, I found that we were to run a command dscSwSys with an argument v15_5, "...then reboot all of the VxWorks ICC boards (if the dscSwSys script didn't do that already)." The script ran and looked like it was attempting to restart the boards but didn't actually do it. Before it would run it complained that a "ps -ef" showed all kinds of active processes that would have to end before the script could run (mostly iop.bin's, including the servers). I quit or kill -9'ed all those processes manually then successfully ran the script, but had to reboot the crates manually anyway. After the reboot all seems well; starting an *OP gives 15_5 as the current ICC version. Spectro Focus ------------- Motors before focus check: sp1 Coll_motor_A -524 Coll_motor_B -525 Coll_motor_C -24 sp2 Coll_motor_A 6118 Coll_motor_B 3068 Coll_motor_C 3618 Temperatures at time of check: tpm_TM_PMSSAvg 7.72071 tpm_TM_TelairAvg.VALA 0.44 Tonight's forecast calls for a low temperature around 0 deg C. Current airTemp is about 11 deg C. I'll focus for 5 deg C. goSpecFocus gives: sp1 red l-r = 0.648 pixels , tolerance = 0.15 pixels, not focused yet Predict a movement of -5929 steps for sp1 red Issue: sp1; mechPiston -5929 Calculating pixel shift maps now... Done calculating pixel shift maps... sp2 red l-r = 0.707 pixels , tolerance = 0.15 pixels, not focused yet Predict a movement of -6469 steps for sp2 red Issue: sp2; mechPiston -6469 Calculating pixel shift maps now... Done calculating pixel shift maps... sp1 blue l-r = -0.420 pixels , tolerance = 0.15 pixels, not focused yet Predict a movement of 15 steps for sp1 blue If red is in focus then you need to rotate blue camera ring by 15 degrees Calculating pixel shift maps now... Done calculating pixel shift maps... sp2 blue l-r = -0.865 pixels , tolerance = 0.15 pixels, not focused yet Predict a movement of 31 steps for sp2 blue If red is in focus then you need to rotate blue camera ring by 31 degrees collimate gives: Reading /data/spectro/53656/sdR-b1-00034237.fit Reading /data/spectro/53656/sdR-b1-00034238.fit COLLIMATE: Min offset = -1.17 pix COLLIMATE: Max offset = 0.15 pix COLLIMATE: RMS across CCD = 0.37 pix COLLIMATE: Mean offset = -0.42 pix COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of 15.0 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-53656-b1-00034237.log PostScript file = Collimate-53656-b1-00034237.ps Reading /data/spectro/53656/sdR-b2-00034237.fit Reading /data/spectro/53656/sdR-b2-00034238.fit COLLIMATE: Min offset = -1.25 pix COLLIMATE: Max offset = -0.26 pix COLLIMATE: RMS across CCD = 0.26 pix COLLIMATE: Mean offset = -0.79 pix COLLIMATE: Camera b2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of 28.2 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-53656-b2-00034237.log PostScript file = Collimate-53656-b2-00034237.ps Reading /data/spectro/53656/sdR-r1-00034237.fit Reading /data/spectro/53656/sdR-r1-00034238.fit COLLIMATE: Min offset = 0.08 pix COLLIMATE: Max offset = 1.23 pix COLLIMATE: RMS across CCD = 0.33 pix COLLIMATE: Mean offset = 0.60 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of -5515 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston -5515" Log file = Collimate-53656-r1-00034237.log PostScript file = Collimate-53656-r1-00034237.ps Reading /data/spectro/53656/sdR-r2-00034237.fit Reading /data/spectro/53656/sdR-r2-00034238.fit COLLIMATE: Min offset = 0.07 pix COLLIMATE: Max offset = 0.99 pix COLLIMATE: RMS across CCD = 0.24 pix COLLIMATE: Mean offset = 0.64 pix COLLIMATE: Camera r2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of -5829 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston -5829" Issued a "mechPiston 3000" on both cameras to bring them in focus for expected nighttime temperatures. The second iteration of goSpecFocus gives: sp1 red l-r = 0.905 pixels , tolerance = 0.15 pixels, not focused yet Predict a movement of -8281 steps for sp1 red Issue: sp1; mechPiston -8281 Calculating pixel shift maps now... Done calculating pixel shift maps... sp2 red l-r = 0.931 pixels , tolerance = 0.15 pixels, not focused yet Predict a movement of -8519 steps for sp2 red Issue: sp2; mechPiston -8519 Calculating pixel shift maps now... Done calculating pixel shift maps... sp1 blue l-r = -0.631 pixels , tolerance = 0.15 pixels, not focused yet Predict a movement of 23 steps for sp1 blue If red is in focus then you need to rotate blue camera ring by 23 degrees Calculating pixel shift maps now... Done calculating pixel shift maps... sp2 blue l-r = -1.120 pixels , tolerance = 0.15 pixels, not focused yet Predict a movement of 40 steps for sp2 blue If red is in focus then you need to rotate blue camera ring by 40 degrees I gave the wrong sign on the mechPiston move, so I will give it "mechPiston -6000" and call it done.